OpenStudy (anonymous):

Can someone explain how to solve this equation :D 3k(2k+8)(3k-9)=0 A. 4,0,3. B. -9,-3. C. 3,-3. D. -4,0,3. E. 3.

rishavraj (rishavraj):

now u have factorised all terms.. now just put 3k = 0, (2k+8) = 0 and so on

OpenStudy (anonymous):

@robtobey and @rishavraj I need the steps

OpenStudy (anonymous):

so i will know how to solve it

OpenStudy (anonymous):

3k=0, k=0 2k+8=0, k=-4 3k-9=0, k=3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks you guys :D